# WonderWorld: Interactive 3D Scene Generation from a Single Image

### v3

@@ Abstract
We present WonderWorld, a novel framework for interactive 3D scene generation that enables users to interactively specify scene contents and layout and see the created scenes in low latency. The major challenge lies in achieving fast generation of 3D scenes. Existing scene generation approaches fall short of speed as they often require (1) progressively generating many views and depth maps, and (2) time-consuming optimization of the scene geometry representations. We introduce the Fast LAyered Gaussian Surfels (FLAGS) as our scene representation and an algorithm to generate it from a single view. Our approach does not need multiple views, and it leverages a geometry-based initialization that significantly reduces optimization time. Another challenge is generating coherent geometry that allows all scenes to be connected. We introduce the guided depth diffusion that allows partial conditioning of depth estimation. WonderWorld generates connected and diverse 3D scenes in less than 10 seconds on a single A6000 GPU, enabling real-time user interaction and exploration. We demonstrate the potential of WonderWorld for user-driven content creation and exploration in virtual environments. We will release full code and software for reproducibility. Project website: https://kovenyu.com/WonderWorld/.
@@ 1 Introduction
Recently, 3D scene generation has surged in popularity, with many works successfully exploring strong generative image priors and improvements in monocular depth estimation [66, 8, 67, 50]. However, existing 3D scene generation approaches are offline, where the user provides a single starting image or text prompt, and then the system, after tens of minutes to hours, returns a fixed 3D scene or a video of the scene. While offline generation may work for small, isolated scenes or videos, this setup is problematic for many scene generation use cases. For example, in game development, world designers want to iteratively build 3D worlds step-by-step, having control over the generation process and being able to see intermediate generation outcomes with low latency. In VR and video games, users expect new content that is larger and more diverse than the scenes currently generated. In the future, users may desire even more: a system that allows them to freely explore and shape a dynamically evolving, infinite virtual world. All of these motivate the problem of interactive 3D scene generation, where the user can control what and where to generate (or extrapolate) new 3D scenes and see the generated scenes in low latency.
@@
The major bottleneck that prevents interactivity is the low speed of generation. Each generated scene typically requires tens of minutes on two main steps: (1) Progressively generating dense multi-view images and aligning depth maps in order to cover occluded regions [66, 8, 50]. (2) Spending a considerable amount of time optimizing the 3D scene representations to shape appropriate geometry and appearances [70, 20, 17]. Besides speed, another challenge is that the generated scenes have strong geometric distortion along the scene boundary due to misalignment or inaccuracy of estimated depth maps, creating seams among generated scenes.
@@
In this paper, we propose a framework named WonderWorld for interactive scene generation. Our input is a single image that depicts the starting scene, and our output is a set of connected yet diverse 3D scenes. To address the speed issue, our core technique includes a novel scene representation, Fast LAyered Gaussian Surfels (FLAGS), and the algorithm to generate it from a single view. The acceleration comes from two aspects: a single-view layer generation that inpaints occluded regions without progressively generating dense views, and a geometry-based initialization that leverages estimated monocular normal and monocular depth for fast optimization. This allows generating a scene in less than 10 seconds on a single GPU. To mitigate the geometry distortion problem, we introduce a guided depth diffusion method to improve the alignment between the geometry of the newly generated scenes and existing scenes.
@@
WonderWorld unlocks the potential for interactive scene generation, allowing users to extrapolate a single image into vast and immersive 3D scenes. Our approach enables new possibilities for applications in virtual reality, gaming, and creative design, where users can quickly generate and explore diverse 3D scenes. In summary, our contributions are three-folded:
@@
  • We showcase and evaluate interactive generation on various types of scenes, such as nature, city, and campus.
@@ 2 Related Work
Many works on generating novel views from a single image attempted to construct renderable 3D scene representations, such as layered depth images [56, 49], radiance fields [65, 54, 52], multi-plane images [55, 72], and point features [42, 59]. Yet, they only supported generating views within small viewpoint changes w.r.t. the input image, as they only built single static scene representations that do not go beyond the input image. Our FLAGS representation integrates the technical ideas from layered depth images [48] and radiance fields [28], yet we incorporate a generative model to support creating many connected scenes rather than a single one.
Later works explored perpetual view generation that allowed much more significant viewpoint changes. Early examples of scene generation focused on extending a single image into a perpetual video with a given camera trajectory: Infinite Images [25] used image stitching, and Infinite Nature [37] and its follow-up works [35, 6, 5] used image generation models specialized to nature images. Since the advent of generative diffusion models, subsequent work has expanded the scope and domain of this work. BlockFusion [61] generates triplanes to represent expandable terrains. SceneScape [15] generates perpetual scenes from a single prompt. WonderJourney [66] instead uses an LLM to generate diverse content and a point cloud representation for the scenes. WonderJourney is most relevant in that it also aims to generate a sequence of diverse scenes, yet it runs offline and requires tens of minutes to generate a single scene as it requires synthesizing dense views in each scene.
Single 3D scene generation.
Recently, scene generation methods have focused primarily on a single scene, with many explicitly focusing on indoor scenes [10, 2, 20, 33, 21] and cities [36, 62, 63]. Recent methods [71, 43, 11] such as Text2NeRF [70], LucidDreamer [8], and CAT3D [17] generate multi-view images of a scene, and RealmDreamer [50] and DreamScene [34] distill multi-view image and depth to generate a 3D scene. Another line of relevant work focuses on single-image 3D scene reconstruction by explicit pose-conditioning or training on scenes [67, 7, 53, 47]. While these approaches demonstrate improvements in the quality of 3D scene generation, they are offline processes generating a fixed scene that is then provided to the user. Since the scene is fixed, their methods do not allow user interaction, e.g., not enabling the user to choose what and where they want to see. We instead address the problem of interactive 3D scene generation, which requires significant improvements with respect to fast generation and extrapolation.
@@ 3 Approach
We target interactive 3D scene generation. Our goal is to generate a set of diverse yet coherently connected 3D scenes (Figure 1). To this end, we propose WonderWorld, a framework that allows real-time rendering and fast scene generation and extrapolation.
@@
We show an illustration of our WonderWorld framework in Figure 2. We start by generating a 3D scene from an input image. Then, the outer control loop keeps iterating over two main steps: generating a scene image and generating FLAGS from the scene image. A user can control where to generate the new scene by moving the camera viewpoint, and control what scene to generate by providing a text prompt. The new scene can be an extrapolation of existing scenes or a standalone scene to be connected later. We summarize the control loop in Alg. 1 in the supplementary material.
@@
The major technical challenge is that we need fast scene generation to allow interactivity. Prior state-of-the-art scene generation methods are slow because they need to progressively generate dense views [66, 8, 50, 20, 70] and spend a long time optimizing scene geometry (e.g., NeRF [70, 17], mesh [20], and 3DGS [8, 50]). We propose the Fast LAyered Gaussian Surfels (FLAGS) and an algorithm to generate it from a single image. Our approach is fast for two reasons: (1) It removes the need for progressive dense view generation to inpaint occluded contents. Instead, we generate geometric layers from a single view and inpaint occluded contents at the layer level. (2) Our representation design enables fast optimization. In particular, our geometry-based initialization significantly reduces the optimization time of a single layer to $<1$ second. Thus, WonderWorld allows fast scene generation within $10$ seconds per scene and real-time rendering, simultaneously on a single GPU.
Another challenge is the geometric distortion that creates seams when connecting two scenes. To mitigate it, we propose to utilize guided depth diffusion to generate geometry. Guided depth diffusion is robust and flexible, allowing the specification of various geometric constraints.
@@ 3.1 Fast LAyered Gaussian Surfels (FLAGS)
For the scale $\mathbf{s}$, our goal is to find an appropriate initialization that meets two requirements: (1) It should minimize rendering aliasing, that is, it should not lead to high-frequency aliases or holes when slightly changing viewpoints (e.g., moving closer to a scene). (2) It should avoid overly big surfels that cause a lot of screen space overlapping to slow down the optimization. Formally, let the spatial sampling interval of an image (i.e., pixel size) be $1$, then the sampling interval at a surfel is $T_{\text{N}}=d/(f\cos\theta)$ where $\theta$ denotes the angle between the surfel normal $\mathbf{n}$ and the image plane normal $\mathbf{n}_{\text{img}}=[0,0,-1]^{\text{T}}$, and $f$ denotes the focal length (Figure 3). According to the Nyquist sampling theorem, the maximum signal frequency should be $1/(2T_{\text{N}})$. Setting the signal frequency of a surfel to be inverse bandwidth of its Gaussian kernel $1/(2ks_{\text{x}})$, we can solve for the initialization of the scales:
@@ 3.2 Guided Depth Diffusion
The main idea of our guided depth diffusion is to formulate the depth estimation of an extrapolated scene as sampling from a depth distribution conditioned on both the scene image and the partially visible depth, $p(\mathbf{D}_{\text{scene}}\mid\mathbf{I}_{\text{scene}},\mathbf{D}_{\text{guide}},\mathbf{M}_{\text{guide}})$. To this end, we inject the partially visible depth as guidance by modifying the denoiser as follows:
@@ 4 Experiments
In our implementation, we use the Stable Diffusion Inpaint model [46] as our outpainting model. We also use it for inpainting the background layer and sky layer, and for text-to-image generation. We use OneFormer [23] to segment the sky and foreground objects. We estimate normal using the Marigold Normal [27]. We use Marigold Depth [27] as our depth diffusion model. We leave more details in the supplementary material. We will release full code and software upon publication for reproducibility.
@@ 4.2 Ablation study
We perform an ablation study, and evaluate them with the same protocal as in the baseline comparison. We show the quantitative results in Table 4.
@@
We compare with a variant (“w/o geometry”) that removes geometry-based initialization and the surfel design, and instead using 3DGS with MipSplatting [69] with the same estimated depth. We increase the optimization iteration such that it achieves the same PSNR as ours at the generation view. However, this variant fails to synthesize high-quality novel views partly due to alias effects (see Figure 6).
@@
We compare with “w/o layers” which uses only a single layer instead of three. Ours significantly outperforms it in both metrics and human preference, as the layered design in our FLAGS fills occluded regions (Figure 7).
@@
We compare with “w/o guidance”. This variant creates significant seams between generated scenes (Figure 8). Our guided depth diffusion mitigates this issue.
@@
Figure 8: Ablation study on the guided depth diffusion. The two images are rendered at a novel view of a generated scene.
@@ 5 Conclusion
A main limitation of WonderWorld is that the generated scenes only have frontal-facing surfaces so that the view synthesis range is limited to an area around the camera, e.g., we cannot move to the back of an object. Future work might incorporate a 3D object generation module such as GRM [64] to generate individual objects separately from the scene background. A few latest works have demonstrated some success [13, 34] following this pipeline. Another limitation is the difficulty in handling detailed objects, such as trees, leaving “holes” or “floaters” when the viewpoint changes. Therefore, we see WonderWorld as an interactive 3D scene prototyping method, rather than a full end-to-end solution. This invites an exciting future direction: using WonderWorld to interactively prototype a coarse 3D world structure, and then refine scene details, complete objects, and remove floaters with slower but higher-fidelity models such as video diffusion [68].
@@
Appendix A Additional Qualitative Results
We show additional baseline comparison results in Figure 14. We show additional qualitative results in Figure 10,  9,  11. To automate generation, we also use the panoramic camera paths. We use the LLM to generate the scene names.
We show different scenes using the same input image in Figure 12, and different styles in the same virtual world in Figure 13. In Table 5, we show a time analysis of WonderWorld.
Figure 9: Qualitative examples. These examples are generated with scene contents automatically generated by the LLM.
Figure 10: Qualitative examples. These examples are generated with scene contents automatically generated by the LLM.
Figure 11: Qualitative examples. These examples are generated with scene contents automatically generated by the LLM.
Outpainting Layer generation Depth Normal Optim.
2.1s 2.3s 2.5s 0.8s 1.9s
@@ Appendix B Algorithms
4:Runtime user control: Real-time rendering camera pose $\mathbf{C}_{\text{rend}}$, generation camera pose $\mathbf{C}_{\text{gen}}$, (optional) user text prompt $\mathcal{U}$
@@
18:
19:in parallel do
20:       Thread 2: Async generation signal (triggered event)
21:       $\mathbf{C}_{\text{gen}}\leftarrow\mathbf{C}_{\text{rend}}$
22:       $\mathbf{I}_{\text{partial}}\leftarrow\text{Render}(\mathbf{C}_{\text{gen}},\mathcal{G})$triangleright Partial rendered image
23:       $\mathbf{M}\leftarrow\text{Find\_empty\_pixels}(\mathbf{I}_{\text{partial}})$
24:       if $\mathcal{U}$ is empty then
@@
26:             $\mathcal{T}\leftarrow\text{Generate\_by\_LLM}(\mathcal{U})$triangleright Eq. 4
27:       else
28:             $\mathcal{T}\leftarrow\text{Generate\_by\_LLM}(\mathcal{U})$triangleright Eq. 4
29:       end if
30:       $\mathbf{I}_{\text{scene}}\leftarrow\text{Outpaint}(\mathbf{I}_{\text{partial}},\mathbf{M},\mathcal{U})$
31:       $\mathcal{G}\leftarrow\text{Generate\_FLAGS}(\mathbf{I}_{\text{scene}},\mathbf{M},\mathcal{T},\mathcal{G})$triangleright Alg. 2
32:end parallel

### v4

@@ Abstract
We present WonderWorld, a novel framework for interactive 3D scene generation that enables users to interactively specify scene contents and layout and see the created scenes in low latency. The major challenge lies in achieving fast generation of 3D scenes. Existing scene generation approaches fall short of speed as they often require (1) progressively generating many views and depth maps, and (2) time-consuming optimization of the scene geometry representations. Our approach does not need to generate multiple views, and it leverages a geometry-based initialization that significantly reduces optimization time. Another challenge is generating coherent geometry that allows all scenes to be connected. We introduce the guided depth diffusion that allows partial conditioning of depth estimation. WonderWorld creates connected and diverse 3D scenes, each generated in less than 10 seconds on a single A6000 GPU, enabling real-time user interaction and exploration. We release full code, software, and interactive demos in https://kovenyu.com/WonderWorld/.
@@ 1 Introduction
Recently, 3D scene generation has surged in popularity, with many works successfully exploring strong generative image priors and improvements in monocular depth estimation [67, 8, 68, 50]. However, existing 3D scene generation approaches are offline, where the user provides a single starting image or text prompt, and then the system, after tens of minutes to hours, returns a fixed 3D scene or a video of the scene. While offline generation may work for small, isolated scenes or videos, this setup is problematic for many scene generation use cases. For example, in game development, world designers want to iteratively build 3D world prototypes step-by-step. This requires having control over the scene contents and layouts while being able to see generation outcomes with low latency. In VR and video games, users expect a world that is larger and more diverse than the scenes currently generated. In the future, users may desire even more: a system that allows them to freely explore and shape a dynamically evolving, infinite virtual world. All of these motivate the problem of interactive 3D scene generation, where the user can control what and where to generate (or extrapolate) a new 3D scene and see how it fits into a world in low latency.
@@
The major bottleneck that prevents interactivity is the low speed of generation. Each generated scene typically requires tens of minutes on two main steps: (1) Progressively generating dense multi-view images and aligning depth maps to cover occluded regions [67, 8, 50]. (2) Spending a considerable amount of time optimizing the 3D scene representations to shape appropriate geometry and appearances [71, 20, 17]. Besides speed, another challenge is that the generated scenes have strong geometric distortion along the scene boundary due to misalignment or inaccuracy of estimated depth maps, creating seams among generated scenes.
@@
In this work, we propose a framework named WonderWorld for interactive scene generation. Our input is a single image that depicts the starting scene, as well as online user controls of camera movement and content prompts. Our output is a set of coherently connected 3D scenes, forming a comprehensive world, according to the online user controls. To address the speed issue, our core technique includes a novel scene representation, Fast LAyered Gaussian Surfels (FLAGS), and the algorithm to generate it from a single view. This allows generating a scene (i.e., the visual and geometric content conditioned on a text prompt and any existing scenes) in less than 10 seconds on a single GPU. To mitigate the geometry distortion problem, we introduce a guided depth diffusion method to improve the alignment between the geometry of the newly generated scenes and existing scenes.
@@
WonderWorld unlocks the potential for interactive scene generation, allowing users to extrapolate a single image into a vast and immersive 3D world. Our approach enables new possibilities for applications in virtual reality, gaming, and creative design, where users can quickly generate and explore diverse 3D worlds. In summary, our contributions are three-folded:
@@
  • We showcase and evaluate interactive generation on various examples, such as nature, city, and campus.
@@ 2 Related Work
Many works on generating novel views from a single image attempted to construct renderable 3D scene representations, such as layered depth images [56, 49], radiance fields [66, 54, 52], multi-plane images [55, 73], and point features [42, 59]. Yet, they only supported generating views within small viewpoint changes w.r.t. the input image, as they only built single static scene representations that do not go beyond the input image. Our FLAGS representation integrates the technical ideas from layered representations [48, 73] and radiance fields [28], yet we focus on a generative task to support creating many connected scenes rather than a single one.
3D world generation.
Later works explored generating more significant viewpoint changes and potentially multiple connected scenes. Early examples of extended scene generation focused on extending a single image into a perpetual video with a given camera trajectory: Infinite Images [25] used image stitching, and Infinite Nature [37] and its follow-up works [35, 6, 5] used image generation models specialized to nature images. Since the advent of generative diffusion models, subsequent work has expanded the scope and domain of this work. BlockFusion [61] generates triplanes to represent expandable terrains. SceneScape [15] generates perpetual scenes from a single prompt. WonderJourney [67] instead uses an LLM to generate diverse content and a point cloud representation for the scenes. WonderJourney is most relevant in that it also aims to generate a sequence of diverse scenes, yet it runs offline and requires tens of minutes to generate a single scene as it requires synthesizing dense views in each scene. Another line of work in large-scale world generation focuses almost entirely on cities [36, 62, 63], producing large-scale 3DGS representations.
3D scene generation.
Recently, scene generation methods have focused primarily on a single, local 3D area, with many explicitly focusing on indoor scenes [10, 2, 20, 33, 21]. Recent methods [72, 43, 11, 65] such as Text2NeRF [71], LucidDreamer [8], and CAT3D [17] generate multi-view images of a scene, and RealmDreamer [50] and DreamScene [34] distill multi-view image and depth to generate a 3D scene. Another line of relevant work focuses on single-image 3D scene reconstruction by explicit pose-conditioning or training on scenes [68, 7, 53, 47]. While these approaches demonstrate improvements in the quality of 3D scene generation, they are offline processes generating a fixed scene that is then provided to the user. Since the scene is fixed, their methods do not allow user interaction, e.g., not enabling the user to choose what and where they want to see. We instead address the problem of interactive 3D scene generation, which requires significant improvements for fast generation and extrapolation.
@@ 3 Approach
We target interactive 3D scene generation. Our goal is to generate a set of diverse yet coherently connected 3D scenes $\{\mathcal{E}_{0},\mathcal{E}_{1},\ldots\}$ from an initial image $\mathbf{I}_{0}$, as well as runtime user controls of camera movements $\mathbf{C}_{\text{gen}}$ and text prompt $\mathcal{U}$ for each scene (Figure 1). Note that we define a single scene, $\mathcal{E}_{i}$, as the visual and geometric content of a text prompt, designed to be consistent with the prior scenes. To this end, we propose WonderWorld, a framework that allows real-time rendering and fast scene generation and extrapolation.
@@
We show an illustration of our WonderWorld framework in Figure 2. We start by generating a 3D scene from an input image. Then, the outer control loop keeps iterating over two main steps: generating a scene image and generating FLAGS from the scene image. A user can control where to generate a new scene by moving the camera, and control the contents by providing a prompt. The new scene can be an extrapolation of existing scenes or a standalone scene to be connected later. We summarize the control loop in Alg. 1 in the supplementary material.
@@
The major technical challenge is that we need fast scene generation to allow interactivity. Prior scene generation methods are slow because they need to progressively generate dense views [67, 8, 50, 20, 71] and spend a long time optimizing scene geometry (e.g., NeRF [71, 17], mesh [20], and 3DGS [8, 50]). We propose the Fast LAyered Gaussian Surfels (FLAGS, Sec. 3.1) and an algorithm to generate it from a single image. Our approach is fast for two reasons. First, it removes the need for progressive dense view generation to inpaint occluded contents. Instead, we generate geometric layers from a single view and inpaint occluded contents at the layer level. Second, our representation design enables fast optimization. In particular, our geometry-based initialization significantly reduces the optimization time of a single layer to $<1$ second. Thus, WonderWorld allows fast scene generation within $10$ seconds per scene and real-time rendering, simultaneously on a single GPU.
Another challenge is the geometric distortion that creates seams when connecting two scenes. To mitigate it, we propose to utilize the guided depth diffusion to generate geometry (Sec. 3.2).
@@ 3.1 Fast LAyered Gaussian Surfels (FLAGS)
For the scale $\mathbf{s}$, our goal is to find an appropriate initialization that meets two requirements: (1) It should minimize rendering aliasing; that is, it should not be too small, which would cause holes when slightly changing viewpoints (e.g., moving closer to a scene). (2) It should avoid overly big surfels that cause a lot of screen space overlapping to slow down the optimization. Formally, let the spatial sampling interval of an image (i.e., pixel size) be $1$, then the sampling interval at a surfel is $T_{\text{N}}=d/(f\cos\theta)$ where $\theta$ denotes the angle between the surfel normal $\mathbf{n}$ and the image plane normal $\mathbf{n}_{\text{img}}=[0,0,-1]^{\text{T}}$, and $f$ denotes the focal length (Figure 3). According to the Nyquist sampling theorem, the maximum signal frequency should be $1/(2T_{\text{N}})$. Setting the signal frequency of a surfel to be inverse bandwidth of its Gaussian kernel $1/(2ks_{\text{x}})$, we can solve for the initialization of the scales:
@@ 3.2 Guided Depth Diffusion
The main idea of our guided depth diffusion is to formulate the depth estimation of an extrapolated scene as sampling from a depth distribution conditioned on both the scene image and the partially visible depth, $p(\mathbf{D}_{\text{scene}}\mid\mathbf{I}_{\text{scene}},\mathbf{D}_{\text{guide}},\mathbf{M}_{\text{guide}})$. To this end, we inject the partially visible depth as guidance by modifying the denoiser as
@@ 4 Experiments
In our implementation, we use the Stable Diffusion Inpaint model [46] as our outpainting model. We also use it for inpainting the background layer and sky layer, and for text-to-image generation. We use OneFormer [23] to segment the sky and foreground objects. We estimate normal using the Marigold Normal [27]. We use Marigold Depth [27] as our depth diffusion model. We leave more details in the supplementary material. We have released full code and software for reproducibility.
@@ 4.2 Ablation study
We perform ablation studies using the same protocol as the baseline comparison, with quantitative results in Table 4.
@@
We compare our model with a variant (“w/o geometry”) that removes geometry-based initialization and the surfel design, and instead uses 3DGS with MipSplatting [70] based on the same estimated depth. We increase the optimization iteration such that it achieves the same PSNR as ours at the generation view. However, this variant fails to synthesize high-quality novel views partly due to alias effects (see Figure 6).
@@
We compare our model with “w/o layers”, which uses only a single layer instead of three. Ours significantly outperforms it in both metrics and human preference, as the layered design in our FLAGS fills occluded regions (Figure 7).
@@
We compare our model with “w/o guidance”. This variant creates significant seams between generated scenes (Figure 8). Our guided depth diffusion mitigates this issue. We show depth alignment evaluation in the supplementary material.
@@
Figure 8: Ablation study on the guided depth diffusion. The two images are rendered with a novel view of a generated scene.
@@ 5 Conclusion
A limitation is that the generated scenes only have frontal-facing surfaces, so the view synthesis range is limited to an area around the camera, as the back side of the object is not generated. Future work may incorporate a 3D object generation module such as GRM [64] to generate individual objects separately from the scene background. A few latest works have demonstrated some success [13, 34] following this pipeline. Another limitation is the difficulty in modeling detailed objects, such as trees, which leave “holes” or “floaters” when the viewpoint changes. Therefore, we see WonderWorld as an interactive 3D world prototyping method, rather than a full end-to-end solution. This invites an exciting future direction: using WonderWorld to interactively prototype a coarse 3D world structure, and then refine scene details and complete objects with slower but higher-fidelity models such as video diffusion [69].
Acknowledgments.
This work is in part supported by NSF RI #2211258, NSF CIF 1955864, NSF PHY-2019786 (http://iaifi.org/), ONR YIP N00014-24-1-2117, N00014-23-1-2355, MURI N00014-22-1-2740, Air Force Artificial Intelligence Accelerator under FA8750-19-2-1000, the Stanford Human-Centered Institute (HAI), Google, Quanta Computer. We thank Yue Gao for proofreading and helping with experiments.
@@ Acknowledgments.
Appendix A Overview
In this supplementary material, we show the following contents:
  • Algorithms of WonderWorld (B)
  • Details on guided depth diffusion (C)
  • Further implementation details (D)
  • Additional experiment results (E)
We also compile video results and interactive viewing examples of the generated virtual worlds in https://kovenyu.com/WonderWorld/. We strongly encourage the reader to view the project website.
@@ Appendix B Algorithms
4:Runtime user control: Real-time camera pose $\mathbf{C}_{\text{rend}}$, generation camera pose $\mathbf{C}_{\text{gen}}$, (optional) user text prompt $\mathcal{U}$
@@
18:in parallel do
19:       Thread 2: Async generation signal (triggered event)
20:       $\mathbf{C}_{\text{gen}}\leftarrow\mathbf{C}_{\text{rend}}$
21:       $\mathbf{I}_{\text{partial}}\leftarrow\text{Render}(\mathbf{C}_{\text{gen}},\mathcal{G})$triangleright Partial rendered image
22:       $\mathbf{M}\leftarrow\text{Find\_empty\_pixels}(\mathbf{I}_{\text{partial}})$
23:       if $\mathcal{U}$ is empty then
@@
25:             $\mathcal{T}\leftarrow\text{Generate\_by\_LLM}(\mathcal{U})$triangleright Eq. 4
26:       else
27:             $\mathcal{T}\leftarrow\text{Generate\_by\_LLM}(\mathcal{U})$triangleright Eq. 4
28:       end if
29:       $\mathbf{I}_{\text{scene}}\leftarrow\text{Outpaint}(\mathbf{I}_{\text{partial}},\mathbf{M},\mathcal{U})$
30:       $\mathcal{G}\leftarrow\text{Generate\_FLAGS}(\mathbf{I}_{\text{scene}},\mathbf{M},\mathcal{T},\mathcal{G})$triangleright Alg. 2
31:end parallel